I am the dad here and frustrated I can't help my son with this problem. It reads: If 1.0 km cubed of air has mass 1,200,000,000 kg, what is the density of air in g/cm cubed? The answer is given as 1.2x10 to the negative 3 g/cm cubed. We know the equation is density equals mass/volume but we obviously are missing something in the "power of" and the ending up with a negative power in the answer. I think we got the conversions correct, but am not even sure of that now. My son and I have worked on this (what appears to be a) simple math problem for over an hour and cannot explain the steps on getting to the answer given. I'm ashamed to admit we're both snowed. Any help would be appreciated. This is from a worksheet and I have no idea if it is from a textbook.: I am the dad here and frustrated I can't help my son with this problem. It reads: If 1.0 km cubed of air has mass 1,200,000,000 kg, what is the density of air in g/cm cubed? The answer is given as 1.2x10 to the negative 3 g/cm cubed. We know the equation is density equals mass/volume but we obviously are missing something in the "power of" and the ending up with a negative power in the answer. I think we got the conversions correct, but am not even sure of that now. My son and I have worked on this (what appears to be a) simple math problem for over an hour and cannot explain the steps on getting to the answer given. I'm ashamed to admit we're both snowed. Any help would be appreciated. This is from a worksheet and I have no idea if it is from a textbook.
Answer by jim_thompson5910(9368)

Remember,

1 kg = 1000 g or 1 kg = 1*10^3 g

and

1 km = 1000 m = 100,000 cm or 1 km = 1*10^3 m = 1*10^5 cm


So 1,200,000,000 kg = 1,200,000,000,000 g or 1,200,000,000 kg = 1.2*10^12 g


and 1 km^3 = (100,000)*(100,000)*(100,000) cm^3 = 1,000,000,000,000,000 cm^3 = 1*10^15 cm^3


-----------------------------

So

1,200,000,000 kg = 1.2*10^12 g

and

1 km^3 = 1*10^15 cm^3


----------------------------------

Also, remember that

 
            Mass
Density =  ---------
            Volume



which is symbolically

D=M/V




D=M/V Start with the given equation.



D=(1.2*10^12)/(1*10^15) Plug in the given mass M=1.2*10^12 g and the given volume V=1*10^15 cm^3


D=1.2*10^(12-15) Divide 1.2 by 1 to get 1.2. Divide 10^12 by 10^15 by subtracting


D=1.2*10^(-3) Subtract the exponents.


So the density is 1.2*10^(-3) g/cm^3
Question 153128: I am the dad here and frustrated I can't help my son with this problem. It reads: If 1.0 km cubed of air has mass 1,200,000,000 kg, what is the density of air in g/cm cubed? The answer is given as 1.2x10 to the negative 3 g/cm cubed. We know the equation is density equals mass/volume but we obviously are missing something in the "power of" and the ending up with a negative power in the answer. I think we got the conversions correct, but am not even sure of that now. My son and I have worked on this (what appears to be a) simple math problem for over an hour and cannot explain the steps on getting to the answer given. I'm ashamed to admit we're both snowed. Any help would be appreciated. This is from a worksheet and I have no idea if it is from a textbook.: I am the dad here and frustrated I can't help my son with this problem. It reads: If 1.0 km cubed of air has mass 1,200,000,000 kg, what is the density of air in g/cm cubed? The answer is given as 1.2x10 to the negative 3 g/cm cubed. We know the equation is density equals mass/volume but we obviously are missing something in the "power of" and the ending up with a negative power in the answer. I think we got the conversions correct, but am not even sure of that now. My son and I have worked on this (what appears to be a) simple math problem for over an hour and cannot explain the steps on getting to the answer given. I'm ashamed to admit we're both snowed. Any help would be appreciated. This is from a worksheet and I have no idea if it is from a textbook.
Answer by ankor@dixie-net.com(4531)

I am the dad here and frustrated I can't help my son with this problem. It reads: If 1.0 km cubed of air has mass 1,200,000,000 kg, what is the density of air in g/cm cubed?
:
1,200,000,000 = 1.2(10^9)
:
Convert the mass: 1.2(10^9) kg to grams:
1.2(10^9) * 10^3 = 1.2(10^12) grams
:
To convert 1 cu km to cubic cm
First:
1 km = 10^3 meters = 10^5 cm
then
Find the volume of a cube with sides of 10^5 cm: (10^5)^3 = 10^15 cu/cm
:
Dens = (1.2(10^12))/(10^15) = 1.2(10^((12-15))) = 1.2(10^-3) grams/cu cm
;
Did this help?
Question 153128: I am the dad here and frustrated I can't help my son with this problem. It reads: If 1.0 km cubed of air has mass 1,200,000,000 kg, what is the density of air in g/cm cubed? The answer is given as 1.2x10 to the negative 3 g/cm cubed. We know the equation is density equals mass/volume but we obviously are missing something in the "power of" and the ending up with a negative power in the answer. I think we got the conversions correct, but am not even sure of that now. My son and I have worked on this (what appears to be a) simple math problem for over an hour and cannot explain the steps on getting to the answer given. I'm ashamed to admit we're both snowed. Any help would be appreciated. This is from a worksheet and I have no idea if it is from a textbook.: I am the dad here and frustrated I can't help my son with this problem. It reads: If 1.0 km cubed of air has mass 1,200,000,000 kg, what is the density of air in g/cm cubed? The answer is given as 1.2x10 to the negative 3 g/cm cubed. We know the equation is density equals mass/volume but we obviously are missing something in the "power of" and the ending up with a negative power in the answer. I think we got the conversions correct, but am not even sure of that now. My son and I have worked on this (what appears to be a) simple math problem for over an hour and cannot explain the steps on getting to the answer given. I'm ashamed to admit we're both snowed. Any help would be appreciated. This is from a worksheet and I have no idea if it is from a textbook.
Answer by stanbon(18984)

If 1.0 km cubed of air has mass 1,200,000,000 kg, what is the density of air in g/cm cubed? The answer is given as 1.2x10 to the negative 3 g/cm cubed.
--------------------
Comment: The key is to properly change (km)^3 to (cm)#.
(km)^3 = (1000m)^3 = (1000*100cm)^3 = (10^5cm)^3 = 10^15cm^3
(1.2x10^9 kg/1.0 km^3)(1000g/1kg)[(km)^3/(10^15cm^3)]
----------
Cancel the units to get g/cm^3
Evaluate the quantities as follows:
[1.2x10^9 * 1000 * 1] / [1.0 * 1 * 10^15] = [1.2*10^12]/[10*15] = 1.2x10^-3
-----------
Final answer 1.2x10^-3 g/cm^3
=======================
Cheers,
Stan H.